you are amazingg!! i appreciate all the help thanl you
(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 150 A.C. A sample of 10 Egyptian skull bread 1

(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 4000 B.C. A sample of 6 Egyptian skull bread2
Answer #1

Solution-:

(1) Part -1 By using R-Software we find mean and SD:

> x=c(135,127,138,127,128,141,130,128,132,134);x
[1] 135 127 138 127 128 141 130 128 132 134
> #For mean
> Mean=mean(x);Mean
[1] 132
> #for Standard Deviation
> sd=sd(x);sd
[1] 4.898979

From this output we get,

Skull breadth mean\bar{x}=132 and  Skull breadth standard deviation s=4.8990

(1) Part -2

The 98 % C.I. is gien by;

here, n=10

(\bar{x}- t_{(n-1),\alpha/2}*\frac{s}{\sqrt{n}},\bar{x}+ t_{(n-1),\alpha/2}*\frac{s}{\sqrt{n}})

(132- 2.821*\frac{4.8990}{\sqrt{10}},132+ 2.821*\frac{4.8990}{\sqrt{10}})

t_{(n-1),\alpha/2}=t_{(10-1),0.02/2}=t_{9,0.01}=2.821

By using MS-Excel Function "=TINV(0.02,9)")

(132- 4.37,132+4.37)

(127.63,136.37)

The required C.I. is (127.63,136.37)

(132- 2.821*\frac{4.8990}{\sqrt{10}},132+ 2.821*\frac{4.8990}{\sqrt{10}})

(132- 2.821*\frac{4.8990}{sqrt{(10)}},132+ 2.821*\frac{4.8990}{sqrt{(10)}})

(2) Part -1 By using R-Software we find mean and SD:

> y=c(136,124,137,122,120,119);y
[1] 136 124 137 122 120 119
> #For mean
> Mean=mean(y);Mean
[1] 126.3333
> #for Standard Deviation
> sd=sd(y);sd
[1] 8.066391

From this output we get,

Skull breadth mean\bar{x}=126.33 and  Skull breadth standard deviation s=8.0664

(2) Part -2

The 99 % C.I. is gien by;

here, n=6

(\bar{x}- t_{(n-1),\alpha/2}*\frac{s}{\sqrt{n}},\bar{x}+ t_{(n-1),\alpha/2}*\frac{s}{\sqrt{n}})

(126.33- 4.032*\frac{8.0664}{\sqrt{6}},126.33+ 4.032*\frac{8.0664}{\sqrt{6}})

t_{(n-1),\alpha/2}=t_{(6-1),0.01/2}=t_{5,0.005}=4.032

By using MS-Excel Function "=TINV(0.01,5)")

(126.33- 13.2778,126.33+ 13.2778)

(113.05,139.61)

The required C.I. is (127.63,136.37)

(126.33- 4.032*\frac{8.0664}{\sqrt{6}},126.33+ 4.032*\frac{8.0664}{\sqrt{6}})

(126.33- 4.032*\frac{8.0664}{sqrt{(6)}},126.33+ 4.032*\frac{8.0664}{sqrt{(6)}})